OpenStudy (anonymous):

What are the approximate solutions of 4x2 + 3 = -12x to the nearest hundredth?

OpenStudy (anonymous):

Change it into a quadratic equation: ax^2 + bx + c = 0 To do this Add 12x to both sides. 4x^2 + 12x + 3 = 0 a = 4 b = 12 c = 3 Use the Quadratic formula to find the solutions for x. \[x = \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\] \[x = \frac{ -12 \pm \sqrt{12^2 - 4(4)(3)} }{ 2(4) }\] \[x = \frac{ -12 \pm \sqrt{144 - 48} }{ 8 }\] \[x = \frac{ -12 \pm \sqrt{96} }{ 8 }\] \[x = \frac{ -12 \pm 9.798 }{ 8 }\] \[x = \frac{ -12 + 9.798 }{ 8 }= -0.276\] \[x = \frac{ -12 - 9.798 }{ 8 }= -2.723\] So the solutions are: x = -0.276 and x = -2.723

OpenStudy (anonymous):

hi i need desperate help with this assignment i understand nothing in it and i always get good marks in maths please help me if u can email me to www.nabz@gmail.com please